# Heterogeneous Mixed Traffic Control and Coordination

### v1

@@ Abstract
Urban intersections, filled with a diverse mix of vehicles from small cars to large semi-trailers, present a persistent challenge for traffic control and management. This reality drives our investigation into how robot vehicles (RVs) can transform such heterogeneous traffic flow, particularly at unsignalized intersections where traditional control methods often falter during power failures and emergencies. Using reinforcement learning (RL) and real-world traffic data, we study heterogeneous mixed traffic across complex intersections under gradual automation by varying RV penetration from 10% to 90%. The results are compelling: average waiting times decrease by up to 86% and 91% compared to signalized and unsignalized intersections, respectively. Additionally, we uncover a “rarity advantage,” where less frequent vehicles, such as trucks, benefit the most from RV coordination (by up to 87%). RVs’ presence also leads to lower CO2 emissions and fuel consumption compared to managing traffic via traffic lights. Moreover, space headways decrease across all vehicle types as RV rate increases, indicating better road space utilization. These results highlight the potential of RVs to manage heterogeneous mixed traffic in complex environments with optimized traffic efficiency and minimized environmental impact.
@@ I INTRODUCTION
Intersections play a crucial role in urban traffic systems where vehicles converge and disperse. Traditionally, traffic lights are used as to manage these complex interactions. However, intersections can turn into congestion points, especially when such control methods fail or are absent. Natural disasters, large-scale power failures, or system malfunctions can disable traffic signals, converting typically controlled intersections into unregulated crossings. Additionally, during temporary road configuration due to construction, intersections may lack signalization. These various situations pose significant traffic management challenges for urban areas, forcing them to navigate complex traffic conditions without traffic lights for extended periods [6, 7].
The emergence of robot vehicles (RVs) with varying degrees of self-driving capabilities has the potential to transform urban traffic management. These vehicles can be leveraged to regulate nearby human-driven vehicles (HVs) so that the mixed traffic, consisting of both RVs and HVs, can be optimized for various objectives such as improving efficiency or reducing emissions [8, 9, 10, 11, 12, 13, 14]. When selecting control strategies for mixed traffic, reinforcement learning (RL) has demonstrated significant potential due to its model-free characteristics [13, 15, 16, 17]. RL is particularly well-suited for managing multi-agent systems at complex and dynamic environments, e.g., traffic at unsignalized intersections, because of its flexibility to adapt and learn effective strategies [18]. Unlike traditional rule-based systems, RL can potentially discover novel solutions that might not be apparent to human traffic engineers [19]. This combination of RVs and RL presents a promising approach for addressing the challenges of mixed traffic control and coordination.
Despite the aforementioned advancements, a significant gap exists between current research of mixed traffic and the complexities of real-world traffic: nearly all existing studies assume homogeneous traffic. Real-world traffic, on the other hand, features a variety of vehicle types, i.e., heterogeneous traffic, equipped with a wide range of kinematic and dynamic parameters. These factors are essential for effective traffic control and management [1, 2, 3, 4]. The limitations of homogeneous traffic become more pronounced when factoring in the complexity of test environments and varying RV penetration rates, raising several important research questions.
  • Is it feasible to implement heterogeneous mixed traffic control in complex environments?
  • What are the performance implications for entire traffic and different vehicle types in terms of efficiency and environmental impact?
  • How do varying RV penetration rates affect performance outcomes?
@@ II RELATED WORK
Several studies explore the management of traffic at unsignalized intersections, particularly for connected and autonomous vehicles (CAVs). Early work by Dresner and Stone [24] introduces a multi-agent intersection control mechanism where CAVs reserve space-time slots in a first-come, first-serve (FCFS) manner. Jin et al. [25] expand on this pioneer work by applying FCFS scheduling to CAV platoons, while Chen et al. [26] propose a controllable gap strategy that adjusts the time gap between vehicles based on their speed and conflict relationships to prevent collisions. Researchers also examine decentralized approaches, such as the consensus-based control method for managing the trajectories of CAVs by Mirheli et al. [27] and Malikopoulos et al.’s work [28] for optimizing the energy consumption. These methods typically assume a fully autonomous environment.
Recently, studies show the effectiveness of RL in managing and optimizing mixed traffic across various scenarios. Wang et al. [17, 29] demonstrate the effectiveness of decentralized multi-agent RL approach for managing mixed traffic at complex, unsignalized intersections, showing significant reductions in waiting time with high RV penetration. Villarreal et al. [10] explore the use of image observations instead of precise traffic information for mixed traffic control via RL. Poudel et al. [30, 31] propose an RL framework that incorporates real-world driving profiles into mixed traffic control and enhances traffic safety, stability, and efficiency. Yan and Wu [32] introduce a model-free RL method for controlling mixed traffic at unsignalized intersections using multi-agent policy decomposition. Peng et al. [33] apply RL to improve mixed traffic efficiency at unsignalized intersections via CAVs. Similarly, Shi et al. [34] present how RL-based method can effectively reduce collisions and improve traffic flow at urban unsignalized intersections compared to traditional control methods.
@@ III-A RL-based Heterogeneous Traffic Control
We formulate the problem of managing heterogeneous mixed traffic at unsignalized intersections as a Partially Observable Markov Decision Process (POMDP), defined as the collection $(\mathcal{X},\mathcal{U},\mathcal{P},\mathcal{R},\mathcal{Y},\mathcal{V},\mathcal{P}_{0},\gamma)$. Here, $\mathcal{X}$ represents the set of all possible states; $\mathcal{U}$ denotes the set of available actions; $\mathcal{P}({x}^{\prime}|{x},u)$ describes the probability of transitioning to a new state ${x}^{\prime}$ given the present state ${x}$ and action $u$; and $\mathcal{R}({x},u)$ defines the reward function that provides feedback based on the state-action pair. The observation space $\mathcal{Y}$ captures the partial view of the environment each RV receives, with $\mathcal{V}(y|{x})$ representing the probability of receiving an observation $y$ from the state ${x}$. $\mathcal{P}_{0}$ is the distribution of initial states, and $\gamma$ represents the factor for discounting future rewards. At each step $t$, an RV chooses an action $u_{t}\in\mathcal{U}$ using the policy $\pi_{\theta}(u_{t}|{x}_{t})$. The environment transitions to the next state ${x}_{t+1}$, and the RV receives a reward $r_{t}$. The goal is to maximize the total discounted reward: $\mathcal{R}_{t}=\sum_{i=t}^{T}\gamma^{i-t}r_{i}$ [17].
RVs decide between two possible actions: Stop and Go. Stop halts the RV before entering the intersection, while Go allows it to proceed. The observation $y_{t}$ of an RV at time $t$ encompasses traffic conditions at the intersection:
$y_{t}=\bigoplus_{k\in K}\langle l_{k},w_{k}\rangle\bigoplus_{k\in K}o_{k},$ (1)
where $K$ is the set of traffic directions, $l_{k}$ is the queue length in direction $k$, $w_{k}$ is the average waiting time in direction $k$, and $o_{k}$ shows the intersection occupancy status.
@@
We design the reward function to encourage traffic efficiency (reduced average waiting times) and discourage vehicle movement conflicts. At time step $t$, the reward $r_{t}$ is defined as $r_{t}=\alpha h_{t}+c_{t}$, where $h_{t}$ is the local reward, $\alpha$ is the weight for local rewards, and $c_{t}$ is the conflict penalty. $h_{t}$ is defined as:
$h_{t}=\begin{cases}-w_{k},&\text{if Stop};\\ w_{k},&\text{if Go},\end{cases}$
and $c_{t}$ is defined as:
$c_{t}=\begin{cases}-1,&\text{if conflict occurs};\\ 0,&\text{otherwise}.\end{cases}$
III-B Heterogeneous Traffic and Complex Intersections
@@
To construct heterogeneous traffic, we analyze the real-world traffic dataset I-24 MOTION [35] and extract the vehicle types and their ratios within: passenger cars (70%), pickups (3%), vans (15%), semi-trailers (11%), and trucks (1%) [31]. The kinematic and dynamic parameters of these vehicle types are presented in Table I [35, 36]. Our study scenarios are four real-world intersections shown in Fig. 1, with the number of incoming lanes 21, 19, 18, and 16 for intersection I, II, III, and IV, respectively.
@@
III-C Evaluation Metrics
We provide comprehensive analysis of heterogeneous mixed traffic in terms of traffic efficiency, environmental impact, and vehicle distribution using the following metrics.
III-C1 Average Waiting Time ($W_{\text{avg}}$)
Average waiting time is defined as the mean time that vehicles spend waiting at the intersection [17]: $W_{\text{avg}}=\frac{1}{N}\sum_{i=1}^{N}W_{i}$, where $N$ is the total number of vehicles and $W_{i}$ is the waiting time of vehicle $i$. We use this metric to assess the overall traffic efficiency. Lower values indicate smoother traffic flow and reduced congestion.
III-C2 CO2 Emission ($E_{\text{CO}_{2}}$)
[…]
III-C3 Fuel Consumption ($F_{\text{avg}}$)
Fuel consumption is the average amount of fuel consumed by vehicles within the intersection over a specified time period [37]: $F_{\text{avg}}=\frac{1}{N}\sum_{i=1}^{N}F_{i}$. $N$ is the total number of vehicles and $F_{i}$ is the fuel consumed by vehicle $i$. This metric provides insights into the energy efficiency of the traffic system.
III-C4 Space Headway ($H_{\text{space}}$)
@@ IV-A Experiment Setup
We model RVs as passenger cars to streamline the complexity of the simulations while still capturing the essence of mixed traffic dynamics. Passenger cars represent the most common vehicle type on urban roads, making them a practical choice for evaluating RV control and coordination. To evaluate the performance of heterogeneous mixed traffic, we compare two baselines: (1) HV-Signalized: all HVs with traffic signals, representing the traditional intersection management method; and (2) HV-Unsignalized: all HVs without traffic signals, representing the unsignalized scenarios. The selected metrics—average waiting time, CO2 emissions, fuel consumption, and space headway—enable us to holistically assess traffic efficiency and environmental impact.
For training the RL policy, we use Rainbow DQN [39] with 1,000 iterations, leveraging an Intel i9-13900KF processor and NVIDIA GeForce RTX 4090 graphics card. The policy’s architecture is a three-layer neural network with 512 units per hidden layer. The learning process uses a rate of 0.0005 and a 0.99 discount factor. RV penetration rates vary from 10% to 90%, incrementing by 10%. We perform 10 evaluations for each RL policy, averaging the outcomes over a 1000-second timeframe. The vehicle type ratios are adjusted based on the RV rates as shown in Table II.
@@
RV (%) Passenger Car (non-RV) (%) Pickup (%) Van (%) Semi-trailer (%) Truck (%)
10 60 3 15 11 1
20 50 3 15 11 1
30 40 3 15 11 1
40 30 3 15 11 1
50 20 3 15 11 1
60 10 3 15 11 1
70 0 3 15 11 1
80 0 2 10 7.3 0.7
90 0 1 5 3.7 0.3
@@ IV-B Overall Traffic Efficiency
The traffic demands for the four unsignalized intersections are extracted from real-world data [17]. Specifically, intersection I has 1,157 vehicles per hour (v/h), intersection II has 1,089 v/h, intersection III sees 928 v/h, and intersection IV experiences the lowest demand at 789 v/h.
TABLE III: Average Waiting Time (s) at four unsignalized intersections under different traffic control scenarios. Increasing RV rates lead to significant reductions in waiting times across all intersections compared to HV-Signalized and HV-Unsignalized, especially when RV penetration rate is 60% or above. For intersection I, waiting time reduces from 122.35s at 10% RV rate to 12.90s at 90% RV rate. Intersection II achieves its best performance at 80% RV rate, with an 86.39% reduction compared to HV-Signalized. Intersections III and IV demonstrate more variable improvements, but still achieve significant reductions in waiting times at higher RV rates (e.g., 60% and 50%, respectively).
@@
Table III highlights the results for average waiting time at four intersections under different scenarios. Across all four intersections, increasing the RV rate leads to substantial reductions in average waiting times compared to both baseline scenarios, indicating improved traffic efficiency. At intersection I, the waiting time decreases from 122.35 s at 10% RV to 12.90 s at 90% RV, achieving maximum improvements of 65.48% and 89.92% over HV-Signalized and HV-Unsignalized. At intersection II, the highest improvement comes at RV rate 80% (8.32 s), providing a 86.39% gain over HV-Signalized and 67.42% over HV-Unsignalized baselines. Intersections III and IV also demonstrate significant gains, with maximum improvements of 50.13% and 77.54% over signalized, and 77.65% and 91.19% over unsignalized baselines, respectively. These results highlight the potential of RVs in dramatically reducing congestion and improving throughput of mixed traffic, compared to traditional control method via traffic lights. Fig. 2 shows an example scenario at intersection I.
Figure 2: Traffic conditions at intersection I at 10-minute mark, RVs in red and HVs in white. (a) HV-Signalized: Traffic flow is controlled via traffic lights. (b) HV-Unsignalized: Intersection quickly becomes congested without traffic lights, resulting in severe gridlock in all directions. (c) Mixed traffic at 10% RV rate: While slightly better than HV-Unsignalized, HVs still dominate the flow, limiting the RVs’ ability to fully optimize the intersection. (d) Mixed traffic at 60% RV rate: RVs learn to adapt to unsignalized conditions, significantly reducing congestion and improving traffic flow. This highlights the potential of RV-based traffic management in complex intersections.

### v2

@@ Abstract
Urban intersections with diverse vehicle types, from small cars to large semi-trailers, pose significant challenges for traffic control. This study explores how robot vehicles (RVs) can enhance heterogeneous traffic flow, particularly at unsignalized intersections where traditional methods fail during power outages. Using reinforcement learning (RL) and real-world data, we simulate mixed traffic at complex intersections with RV penetration rates ranging from 10% to 90%. Results show that average waiting times drop by up to 86% and 91% compared to signalized and unsignalized intersections, respectively. We observe a “rarity advantage,” where less frequent vehicles benefit the most (up to 87%). Although CO2 emissions and fuel consumption increase with RV penetration, they remain well below those of traditional signalized traffic. Decreased space headways also indicate more efficient road usage. These findings highlight RVs’ potential to improve traffic efficiency and reduce environmental impact in complex, heterogeneous settings.
@@ I INTRODUCTION
Intersections play a crucial role in urban traffic systems where vehicles converge and disperse. Traditionally, traffic lights are used to manage these complex interactions. However, intersections can turn into congestion points, especially when such control methods fail or are absent. Natural disasters, large-scale power failures, or system malfunctions can disable traffic signals, converting typically controlled intersections into unregulated crossings. Additionally, during temporary road configuration due to construction, intersections may lack signalization. These various situations pose significant traffic management challenges for urban areas, forcing them to navigate complex traffic conditions without traffic lights for extended periods [6, 7].
The emergence of robot vehicles (RVs) with varying degrees of self-driving capabilities has the potential to transform urban traffic management. These vehicles can be leveraged to manage nearby human-driven vehicles (HVs) so that the mixed traffic, consisting of both RVs and HVs, can be optimized for various objectives such as improving efficiency or reducing emissions [8, 9, 10, 11, 12, 13]. When selecting control strategies for mixed traffic, reinforcement learning (RL) has demonstrated significant potential due to its model-free characteristics [13, 14, 15]. RL is particularly well-suited for managing multi-agent systems at complex and dynamic environments, e.g., traffic at unsignalized intersections, because of its flexibility to adapt and learn effective strategies [16]. Unlike traditional rule-based systems, RL can potentially discover novel solutions that might not be apparent to human traffic engineers [17]. This combination of RVs and RL presents a promising approach for addressing the challenges of mixed traffic control and coordination.
Despite the aforementioned advancements, a significant gap exists between current research of mixed traffic and the complexities of real-world traffic: nearly all existing studies assume homogeneous traffic. Real-world traffic, on the other hand, features a variety of vehicle types, i.e., heterogeneous traffic, equipped with a wide range of kinematic and dynamic parameters. These factors are essential for effective traffic control and management [1, 2, 3, 4]. The limitations of homogeneous traffic become more pronounced when factoring in the complexity of test environments and varying RV penetration rates, raising several important research questions: 1) Is it feasible to implement heterogeneous mixed traffic control in complex environments? 2) What are the performance implications for entire traffic and different vehicle types in terms of efficiency, environmental impact? and 3) How do varying RV penetration rates affect performance?
@@ II RELATED WORK
Several studies explore the management of traffic at unsignalized intersections, particularly for connected and autonomous vehicles (CAVs). Early work by Dresner and Stone [22] proposes a multi-agent intersection control system where CAVs reserve space-time slots using first-come, first-serve (FCFS). Jin et al. [23] expand on this pioneer work by applying FCFS scheduling to CAV platoons, while Chen et al. [24] propose a controllable gap strategy that adjusts the time gap between vehicles based on their speed and conflict relationships to prevent collisions. Researchers also examine decentralized approaches, such as the consensus-based control method for managing the trajectories of CAVs by Mirheli et al. [25] and Malikopoulos et al.’s work [26] for optimizing the energy consumption. These methods typically assume a fully autonomous environment.
Recently, studies show the effectiveness of RL in managing and optimizing mixed traffic across various scenarios. Wang et al. [15, 27] demonstrate that decentralized multi-agent RL can significantly reduce waiting times at complex unsignalized intersections with high RV penetration. Villarreal et al. [10] explore image-based observations as a substitute for precise traffic data in RL-based control. Poudel et al. [28, 29] propose an RL framework that integrates real-world driving profiles to improve safety, stability, and efficiency. Yan and Wu [30] introduce a model-free multi-agent RL approach for controlling mixed traffic at unsignalized intersections. Peng et al. [31] apply RL to boost efficiency at such intersections using CAVs. Shi et al. [32] show that RL can reduce collisions and improve flow in urban unsignalized settings over traditional control methods.
@@ III-A RL-based Heterogeneous Traffic Control
We formulate the problem of managing heterogeneous mixed traffic at unsignalized intersections as a Partially Observable Markov Decision Process (POMDP), defined as the collection $(\mathcal{X},\mathcal{U},\mathcal{P},\mathcal{R},\mathcal{Y},\mathcal{V},\gamma)$. Here, $\mathcal{X}$ represents the set of all possible states; $\mathcal{U}$ denotes the set of available actions; $\mathcal{P}({x}^{\prime}|{x},u)$ describes the probability of transitioning to a new state ${x}^{\prime}$ given the present state ${x}$ and action $u$; and $\mathcal{R}({x},u)$ defines the reward function that provides feedback based on the state-action pair. The observation space $\mathcal{Y}$ captures the partial view of the environment each RV receives, with $\mathcal{V}(y|{x})$ representing the probability of receiving an observation $y$ from the state ${x}$. The discount factor $\gamma\in[0,1)$ determines the weighting of future rewards. At each step $t$, an RV chooses an action $u_{t}\in\mathcal{U}$ using the policy $\pi_{\theta}(u_{t}|{x}_{t})$. The environment transitions to the next state ${x}_{t+1}$, and the RV receives a reward $r_{t}$. The goal is to maximize the total discounted reward: $\mathcal{R}_{t}=\sum_{i=t}^{T}\gamma^{i-t}r_{i}$ [15].
RVs decide between two possible actions: Stop and Go. Stop halts the RV before entering the intersection, while Go allows it to proceed. This high-level, binary action space allows the policy to focus on strategic coordination, while the precise acceleration and deceleration are handled by low-level controllers described in Sec. III-B. The observation $y_{t}$ of an RV at time $t$ encompasses traffic conditions at the intersection: $y_{t}=\oplus_{k\in K}\langle l_{k},w_{k}\rangle\oplus_{k\in K}\langle o_{k}\rangle\oplus\langle d_{int}\rangle$, where $K$ is the set of traffic directions, $l_{k}$ is the queue length in direction $k$, $w_{k}$ is the average waiting time in direction $k$, $o_{k}$ shows the intersection occupancy status, and $d_{int}$ represents the RV’s distance to the intersection. This observation formulation is consistent with the capabilities of connected vehicles, which can access global traffic states via V2X communication.
@@
We design the reward function to encourage traffic efficiency (reduced average waiting times) and discourage vehicle movement conflicts. When multiple RVs from conflicting streams arrive at the intersection entrance and decide to proceed, priority is assigned based on a combined score of waiting time and queue length. At time step $t$, the reward $r_{t}$ is defined as $r_{t}=\alpha h_{t}+c_{t}$, where $h_{t}$ is the local reward, $\alpha$ is the weight for local rewards, and $c_{t}$ is the conflict penalty. The local reward is given by $h_{t}=-w_{k}$ if Stop, and $h_{t}=w_{k}$ if Go, where $w_{k}$ is the average waiting time of all vehicles moving in direction $k$. The conflict penalty is defined as $c_{t}=-1$ if a conflict occurs, and $c_{t}=0$ otherwise. This reward formulation with explicit penalization of conflicts has been shown to reduce unsafe interactions in similar settings [15].
III-B Vehicle Behavior Implementation
The behavior of HVs is governed by Intelligent Driver Model (IDM) [33], which provides acceleration based on surrounding traffic conditions. RVs use a hybrid approach: they employ IDM when outside an intersection (30m from the intersection) and execute a learned policy when inside an intersection. When the policy indicates forward, the RV accelerates at maximum rate; when signaling to stop, it decelerates based on its current speed $u$ and distance to intersection $d_{int}$: $a=-u^{2}/2d_{int}$.
III-C Complex Intersections and Heterogeneous Traffic
@@
Our study scenarios are four real-world intersections shown in Fig. 1, with the number of incoming lanes 21, 19, 18, and 16 for intersections I, II, III, and IV, respectively. Vehicle routes and trajectories are generated using SUMO’s routing tools based on turning count data [15]. To construct heterogeneous traffic, we analyze the real-world dataset I-24 MOTION [34] and extract the vehicle types and their ratios within: passenger cars (70%), pickups (3%), vans (15%), semi-trailers (11%), and trucks (1%) [29]. We ground the vehicle type distribution of our simulation based on these ratios. The kinematic and dynamic parameters of these vehicle types are presented in Table I [34, 35].
@@
III-D Evaluation Metrics
We provide comprehensive analysis of heterogeneous mixed traffic in terms of traffic efficiency and environmental impact using the following metrics.
III-D1 Average Waiting Time
Average waiting time measures the mean time vehicles spend waiting at the intersection [15], calculated as $W_{\text{avg}}=\frac{1}{N}\sum_{i=1}^{N}W_{i}$, where $N$ is the total number of vehicles and $W_{i}$ is the waiting time of vehicle $i$. This metric assesses traffic efficiency, with lower values indicating smoother flow and less congestion.
III-D2 CO2 Emissions and Fuel Consumption
We evaluate the environmental impact of heterogeneous mixed traffic using CO2 emissions and fuel consumption, estimated with SUMO’s built-in HBEFA3-based model [36, 37]. Each vehicle type is assigned an emission class corresponding to its real-world counterpart (e.g., HBEFA3/PC_G_EU4 for passenger cars, HBEFA3/HDV_D_EU4 for semi-trailers). For each vehicle type $v$, the average CO2 emission is calculated as $E_{\text{CO}_{2}}^{v}=\frac{1}{N_{v}}\sum_{i=1}^{N_{v}}E_{i}$, where $N_{v}$ is the total number of vehicles of type $v$ passing through the intersection, and $E_{i}$ is the CO2 emission from vehicle $i$. We measure fuel consumption in a similar manner. Although the HBEFA3-based model does not fully capture the variability of real-world driving, we use it because of its integration with SUMO and its suitability for comparative analysis across scenarios when empirical data is unavailable.
III-D3 Space Headway
@@ IV-A Experiment Setup
We model RVs as passenger cars to streamline the complexity of the simulations while still capturing the essence of mixed traffic dynamics. Passenger cars represent the most common vehicle type on urban roads, making them a practical choice for evaluating RV control and coordination. Each newly spawned passenger car is randomly designated as an RV or HV based on a pre-specified RV rate.
To evaluate the performance of heterogeneous mixed traffic, we compare two baselines: (1) HV-Signalized: HVs operating under traffic signals, representing the traditional intersection management method; and (2) HV-Unsignalized: HVs without signal control, representing the unsignalized scenarios. Our study is conducted in a heterogeneous traffic environment, where RVs must navigate the complex dynamics introduced by diverse types of HVs. Consequently, we focus on practical, real-world baselines rather than homogeneous ones, which would fail to capture this core challenge.
For each RV penetration rate (10% to 90% in 10% steps), we train a shared RL policy using Rainbow DQN [38] for 1,000 iterations on an Intel i9-13900KF and NVIDIA RTX 4090, with each run taking 20–30 hours. The network has three layers of 512 units, a learning rate of 0.0005, and a discount factor of 0.99. Each policy is evaluated 10 times over 1,000 seconds. Vehicle type ratios follow Table II.
@@
RV (%) Passenger Car (non-RV) (%) Pickup (%) Van (%) Semi-trailer (%) Truck (%) 10 60 3 15 11 1 20 50 3 15 11 1 30 40 3 15 11 1 40 30 3 15 11 1 50 20 3 15 11 1 60 10 3 15 11 1 70 0 3 15 11 1 80 0 2 10 7.3 0.7 90 0 1 5 3.7 0.3
@@ IV-B Overall Traffic Efficiency
The traffic demands for the four unsignalized intersections are extracted from real-world data [15]. Specifically, intersection I has 1,157 vehicles per hour (v/h), intersection II has 1,089 v/h, intersection III sees 928 v/h, and intersection IV experiences the lowest demand at 789 v/h. Table III highlights the results for average waiting time at four intersections under different scenarios. Across all four intersections, increasing the RV rate leads to substantial reductions in average waiting times compared to both baseline scenarios, indicating improved traffic efficiency. At intersection I, the waiting time decreases from 122.35 s at 10% RV to 12.90 s at 90% RV, achieving maximum improvements of 65.48% and 89.92% over HV-Signalized and HV-Unsignalized. At intersection II, the highest improvement comes at RV rate 80% (8.32 s), providing a 86.39% gain over HV-Signalized and 67.42% over HV-Unsignalized baselines. Intersections III and IV also demonstrate significant gains, with maximum improvements of 50.13% and 77.54% over signalized, and 77.65% and 91.19% over unsignalized baselines, respectively. These results highlight the potential of RVs in reducing congestion and improving throughput of mixed traffic, compared to traditional control method via traffic lights. Fig. 2 shows an example scenario at intersection I.
TABLE III: Average Waiting Time (s) at four intersections under different traffic control scenarios. Higher RV rates significantly reduce waiting times across all intersections compared to HV-Signalized and HV-Unsignalized, especially at 60% RV rate or above. For Intersection I, waiting time drops from 122.35s at 10% RV rate to 12.90s at 90%. Intersection II performs best at 80% RV rate, reducing waiting time by 86.39%. Intersections III and IV show variable improvements but still achieve substantial reductions at higher RV rates (e.g., 60% and 50%, respectively).
@@
Figure 2: Traffic at intersection I at the 10-minute mark (RVs in red, HVs in white). (a) HV-Signalized: Flow is controlled by traffic lights. (b) HV-Unsignalized: Severe gridlock emerges without signals. (c) 10% RV: Slight improvement, but HV dominance limits coordination. (d) 60% RV: RVs adaptively coordinate, reducing congestion and enhancing flow.